OpenStudy (erty):

How would I slove ((x^2-9^2) +3^2) + x^2= ?

OpenStudy (erty):

^2 means squared

OpenStudy (erty):

Please help!!!

OpenStudy (astrophysics):

What do you mean by solve?

OpenStudy (astrophysics):

\[(x^2-9^2)+3^2+x^2\]

OpenStudy (astrophysics):

What exactly do you want to do with this

OpenStudy (erty):

I got help from a tutor and he said for me to finish the problem I need to solve that.

OpenStudy (astrophysics):

You mean simplify

OpenStudy (astrophysics):

Just drop the brackets and collect like terms

OpenStudy (erty):

oh and i forgot it equals to 12^2

OpenStudy (astrophysics):

\[(x^2-9^2)+3^2+x^2 = 12^2\]

OpenStudy (erty):

yes

OpenStudy (astrophysics):

Ok first thing, lets drop the brackets and simplify the left side

OpenStudy (erty):

ok

OpenStudy (astrophysics):

So go ahead and try that, and remember 9^2 = 81, and 3^2 = 9

OpenStudy (astrophysics):

12^2 = 144

OpenStudy (astrophysics):

The exponents tell you how many times it's being multiplied by itself

OpenStudy (astrophysics):

9^2 = 9 * 9 = 81

OpenStudy (erty):

I got 6

OpenStudy (astrophysics):

What do you mean

OpenStudy (erty):

I squared them and then got -72

OpenStudy (erty):

then i squared 12

OpenStudy (astrophysics):

just show the math here step by step

OpenStudy (astrophysics):

Because you should get two solutions

OpenStudy (erty):

wait I found a mistake

OpenStudy (erty):

I don't know. I can't get it. The answers I chose from are: A 10.5 B 10 C 6 root 3 D 20

OpenStudy (erty):

Is your answer one of those?

OpenStudy (astrophysics):

\[x^2-81+9+x^2=144\] dropping the brackets and squaring, so simplify the left side a bit what do we get

OpenStudy (astrophysics):

Take it step by step

OpenStudy (erty):

2x^2 -72=144

OpenStudy (astrophysics):

Yes, good now what's next?

OpenStudy (erty):

add 72

OpenStudy (astrophysics):

Right what do you get

OpenStudy (erty):

216

OpenStudy (astrophysics):

Good so we have \[2x^2 = 216\]

OpenStudy (astrophysics):

next?

OpenStudy (erty):

yes

OpenStudy (erty):

divide 2

OpenStudy (astrophysics):

good then we get?

OpenStudy (erty):

108

OpenStudy (erty):

then we square root it

OpenStudy (astrophysics):

Right! When we square we actually get a positive and negative solution but it seems they just want a positive answer so lets say \[x=\sqrt{108}\]

OpenStudy (erty):

but there is no answer like that in my multiple choice

OpenStudy (astrophysics):

Now we need to simplify 108 in the radical

OpenStudy (erty):

the only radical there is, is 6 root 3

OpenStudy (astrophysics):

\[\sqrt{108} = \sqrt{36 \times 3} = 6\sqrt{3}\]

OpenStudy (erty):

ok. Thanks!

OpenStudy (erty):

Thank you! I got it correct!

OpenStudy (astrophysics):

Yw